The weights of 1,000 children, in average, is 49kg with standard deviation of 18kg. Suppose the weights are normally distributed, how many children weigh between 53kg and 59kg?


1
Expert's answer
2022-02-07T16:38:41-0500

Let "X=" weight: "X\sim N(\mu, \sigma^2 )."

Given "\mu=49 \ kg, \sigma=18\ kg."


"P(53<X<59)=P(X<59)-P(X\le 53)"

"=P(Z<\dfrac{59-49}{18})-P(X\le \dfrac{53-49}{18})"

"\approx P(Z<0.555556)-P(Z\leq 0.222222)"

"\approx0.71074264-0.58792955=0.1228"

"0.1228(1000)=123"

123  children weigh between 53kg and 59kg.